# Differential Privacy of Cross-Attention with Provable Guarantee

### v1

@@ G.2 Algorithm Modifications
we can compute $\ell_{2}^{2}$ norm easily (see details in Lemma C.2). We then show how to extend our one dimensional $\ell_{1}$ distance algorithm (Algorithm 6 and 7) to $\ell_{2}^{2}$ distance with minor modifications.
Theorem G.7 (DPTreeDistance $\ell_{2}^{2}$ distance).
@@
The data structure DPTreeDistance (Algorithm 6,7) uses $O(n)$ spaces to solve weighted $\ell_{2}^{2}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
@@
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}$,
@@
The proof is similar to that of Theorem D.6, except that now our additive error includes $R$ increased by a power of $2$, i.e., from $O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$ to $O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$. ∎
@@
Theorem G.8 (Softmax query, formal version of Theorem 4.2).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition G.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
@@
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
@@
The DP proof is the same as the proof of Lemma E.1.
We then show the time complexity. From Lemma G.6, we know that constructing $P(\cdot)$ requires $O(r)$ time. In the first for loop of Init, the dominating time consumption is $O(nr)$. The second for loop also has a time complexity of $O(nr)$. Therefore, the total time complexity for Init is $O(nr)$. In the DistanceQuery function, constructing $P(y)$ takes $O(r)$ time. Within the for loop, it requires $O(\alpha^{-1}r\log^{2}n)$. Thus, the total time complexity for DistanceQuery is $O(\alpha^{-1}r\log^{2}n)$.
@@
The proof of the error follows from the triangle inequality by combining the errors in Lemma G.6 and Theorem G.7.
G.3 Adaptive Softmax
In this section, we show how to make Algorithm 3 robust to adaptive query. We follow the same idea from Section F. We notice that, in the softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$-distance in Section F. Therefore, we need to recalculate Lipschitz constant first.
Lemma G.9 (Lipschitz of weighted softmax).
@@ G.3 Adaptive Softmax
where the first step follows from definition of $Z(y),Z(\widetilde{y})$, the second step follows from triangular inequality, the third step follows from Fact A.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$. ∎
@@
Lemma G.10 (Adaptive softmax, formal version of Lemma 6.8).
@@
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
@@
The proof follows from the same idea as the proof of Lemma F.8, except that we use Theorem G.8 and the Lipschitz in Lemma G.9. ∎
Theorem G.11 (Adaptive query softmax data structure, formal version of Theorem 4.4).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition G.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter, and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnr)$ spaces to solve weighted softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
@@
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
@@
We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma G.10. We know the additive error introduced by adaptive query is $E_{a}:=O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0})$ and the relative error introduced by polynomial kernel approximation is $E_{p}:=w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$. It can be shown that:
@@
G.4 Proof of main result
@@
Theorem G.12 (Cross-attention, formal version of Theorem 5.1).
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma 6.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. Let $\phi$ be softmax function (Definition 4.1). There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
@@
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP,
    • $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
From Section 3.4, we know that we can make matrix $AV$ in cross-attention computation DP. Next, from Theorem 4.4, for $i\in[m],j\in[n],k\in[d]$, we have $(AV)_{i,k}$ is $(\epsilon,\delta+\delta^{\prime})$-DP and also robust to adaptive query. For $D$, we compute the exact true value. By the post-processing property of DP (Fact 3.9), our cross-attention is DP.

### v2

@@ I.2 Algorithm Modifications
we can compute $\ell_{2}^{2}$ norm easily (see details in Lemma E.2). We then show how to extend our one dimensional $\ell_{1}$ distance algorithm (Algorithm 5 and 6) to $\ell_{2}^{2}$ distance with minor modifications.
Theorem I.7 (DPTreeDistance $\ell_{2}^{2}$ distance).
@@
The data structure DPTreeDistance (Algorithm 5,6) uses $O(n)$ spaces to solve weighted $\ell_{2}^{2}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}$,
@@
The proof is similar to that of Theorem F.6, except that now our additive error includes $R$ increased by a power of $2$, i.e., from $O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$ to $O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$. ∎
@@
Theorem I.8 (Softmax query, formal version of Theorem 4.2).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition I.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
@@
The DP proof is the same as the proof of Lemma G.1.
We then show the time complexity. From Lemma I.6, we know that constructing $P(\cdot)$ requires $O(r)$ time. In the first for loop of Init, the dominating time consumption is $O(nr)$. The second for loop also has a time complexity of $O(nr)$. Therefore, the total time complexity for Init is $O(nr)$. In the DistanceQuery function, constructing $P(y)$ takes $O(r)$ time. Within the for loop, it requires $O(\alpha^{-1}r\log^{2}n)$. Thus, the total time complexity for DistanceQuery is $O(\alpha^{-1}r\log^{2}n)$.
@@
The proof of the error follows from the triangle inequality by combining the errors in Lemma I.6 and Theorem I.7. ∎
I.3 Adaptive Softmax
In this section, we show how to make Algorithm 3 robust to adaptive query. We follow the same idea from Section H. We notice that, in the Softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$-distance in Section H. Therefore, we need to recalculate Lipschitz constant first.
Lemma I.9 (Lipschitz of weighted Softmax).
@@ I.3 Adaptive Softmax
where the first step follows from definition of $Z(y),Z(\widetilde{y})$, the second step follows from triangular inequality, the third step follows from Fact B.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$. ∎
@@
Lemma I.10 (Adaptive Softmax, formal version of Lemma C.8).
@@
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
@@
The proof follows from the same idea as the proof of Lemma H.8, except that we use Theorem I.8 and the Lipschitz in Lemma I.9. ∎
Theorem I.11 (Adaptive query Softmax data structure, formal version of Theorem 4.4).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition I.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter, and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnr)$ spaces to solve weighted Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
@@
We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma I.10. We know the additive error introduced by adaptive query is $E_{a}:=O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0})$ and the relative error introduced by polynomial kernel approximation is $E_{p}:=w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$. It can be shown that:
@@
I.4 Proof of Main Result
@@
Theorem I.12 (Softmax cross-attention, formal version of Theorem 3.1).
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma C.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP,
    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
From Section 3, we know that we can ensure matrix $AV$ in cross-attention computation satisfies DP. Next, from Theorem 4.4, for $i\in[m],j\in[n],k\in[d]$, we have $(AV)_{i,k}$ is $(\epsilon,\delta+\delta^{\prime})$-DP and also robust to adaptive query. For $D$, we compute the exact true value. By the post-processing property of DP (Fact B.5), our cross-attention process is DP.